OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (-7,9); m=4/5 a.)y+7=4/5(x-9) b.)y-7=4/5(x+9) c.)y-9=4/5(x+7) d.)y+9=4/5(x-7) Thanks(:

OpenStudy (anonymous):

@itsmylife

OpenStudy (anonymous):

y-9= 4/5(x+7) its y-y1=m(x-x1) ;)

OpenStudy (anonymous):

Ahh you are perfect!(:

OpenStudy (anonymous):

thanks , lady ;)